How many years does it take for an investment to double in value if it is invested at 6%
if interest is compounded quarterly?
if interest is compounded continuously?
thanks

Answer by lwsshak3(11628)

How many years does it take for an investment to double in value if it is invested at 6%
if interest is compounded quarterly?
if interest is compounded continuously?
**
compound interest formula: A=P(1+r)^t, P=initial investment, r=interest rate per period, t=number of periods, A=amount after t periods.
For continuous compounding: A=Pe^rt
..
Quarterly compounding: A/P=2, r=.06/4=.015, t=quarters
A=P(1+r)^t
A/P=(1+.06/4)^t
A/P=(1+.015)^t
2=(1.015)^t
take log of both sides
log(2)=t*log(1.015)
t=log(2)/log(1.015)
t≈46.55 qtrs≈11.64 years
..
Continuous compounding: A/P=2, r=.06, t=years
A=Pe^rt
A/P=e^rt
2=e^rt
2=e^.06t
take log of both sides
ln2=.06t*lne
lne=1
.06t=ln2
t=ln2/.06
t≈11.55 years